OpenStudy (anonymous):

Given F(x)=x-3 and g(x) =1/x^2-9, find [g*f](x)

OpenStudy (anonymous):

A. 1/x^2-6x, x=3 B. 1/x^2-6x, x=0,3,6 C. 28-3x^2/x^2-9, x=3 D. 28-3x^2/x^2-9, x=1,3,6

OpenStudy (anonymous):

...???

OpenStudy (anonymous):

so answer choice C is right

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

oh I'm sorry I think it's D

OpenStudy (anonymous):

i'm sorry i think i'm doing this wrong

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the right answer is C

OpenStudy (anonymous):

uhg i keep putting it wrong -.-

OpenStudy (anonymous):

\[\frac{ 1 }{ (x-3)^2-9 }\] THERE WE GO, I'm just distracted right now

OpenStudy (anonymous):

ok so can you please tell me which one is the right answer from this answer choices a the top

OpenStudy (anonymous):

Lets figure it out

OpenStudy (anonymous):

\[\frac{ 1 }{ (x-3)^2-3^2 } \implies \frac{ 1 }{ (x-3-3)(-3+x+3) }\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 1 }{ x(x-3-3) }\] after combining like terms we get this, check me on my math :P \[\implies \frac{ 1 }{ x(x-6) }\]

OpenStudy (anonymous):

So that does equal \[\frac{ 1 }{ x^2-6x }\]

OpenStudy (anonymous):

I was trying to figure if its a or b because I'm not sure you puy x\[x \neq 3 or x \neq 0,3, 6\]

OpenStudy (anonymous):

A lot of them don't really make sense huh?

OpenStudy (anonymous):

ya that's what I'm confuse with.....

OpenStudy (anonymous):

Ok give me one second lol, I'll get back to this.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok so my first thought was right, (g*f)(x) is just multiply bleh!

OpenStudy (anonymous):

Is the question really asking (g*f)(x) because that is just multiply... @nehapa

OpenStudy (anonymous):

it's asking for